In a mechanism as shown in Fig. 8.39, the link AB rotates with a uniform angular velocity of 30 rad/s. The lengths of various links are : A B = 100 mm ; BC = 300 mm ; BD = 150 mm ; DE = 250 mm ; EF = 200 mm ; DG = 165 mm.

Determine the velocity and acceleration of G for the given configuration.



1
Expert's answer
2021-06-07T05:45:52-0400


ad"=3.5 cm = \frac{3.5}{1.667}= 2.09m/sec=V_D"

"ag=1cm = \frac{1}{1.667}=0.59m/sec=V_G \space velocity \space of \space G =0.59m/sec"

"ae=0.5cm=\frac{0.5}{1.667}=0.29m/sec =V_E"

"bc=4.5cm=2.699m\/sec=V_{BC}"

"wbc=\frac{V{BC}}{BC}=\frac{2.699}{0.3}=8.99rad/sec"

"de=3.8=2.27m\/sec=V_{DE}"

"w_{DE}=\frac{V_{DE}}{DE}=\frac{2.27}{0.25}=9.11rad/sec"

"w_{EF}=\frac{VE}{EF}=\frac{0.29}{0.2}=1.45 rad/sec"


"a_{BN}=AB(wAB)^2"

"=0.1(30)^2=90m\/s^2"

"a_{Bcn}=BC(w_{Bc})^2"

"=0.3(8.99)^2=24.24m\/s^2"

"a_{DE}=BE(w_{DE})^2"

"=0.25(9.11)^2=20.74m\/s^2"

"a_{EF}=EF(w{EF})^2"

"0.2(1.45)^2=0.4205m\/s^2"

"(SF)_a=\frac{10}{90}=0.1112cms^2/m"

"ac=18cm=ac=\frac{18}{0.1112}=162.002m/sec^2"

"ad=10cm=ad=\frac{10}{0.112}=90.0009m/s^2"

"ag=19.3cm=ag=\frac{19.3}{0.112}=173.70002m/s^2"

Acceleration of G "=1737m\/sec^2"